find the least natural no having 24 divisors.
(ques from permutation and combination)

Grade:11

1 Answers

moumi roy
7 years ago
the smallest number  may be 360.let Q be the needed number and the total devisors are(a+1)(b+1)..(n+1)=24.as 24 can only be broken inti smaller numbers in only a few waylike 3*2*2*2=4*3*2=6*4=6*2*2=24.so the snallest prime will be out of all 360 as  2^(2-1)*3^(3-1)*5^(2-1)=2^3*3^3*5=360